OpenStudy (anonymous):

A trigonometric question: On an ABC triangle , the edges are a, b and c. (b-c)/(b+c)=1/(root3) m(A)=60 degree angle what is angle of m(B) ?

OpenStudy (amistre64):

if its a right triangle, then the m(B) will equal 30 degrees. if its not a right triangle, then a picture would be helpful

OpenStudy (anonymous):

There is no given shape.Thanks anyway.

OpenStudy (anonymous):

Answer is equal 105 degrees.

OpenStudy (amistre64):

then you could do it the hard way by making a system of equations: b - c = 1 b + c = sqrt(3) and finding your b and c....

OpenStudy (amistre64):

then use the law of cosines to find "a" and the law of sines to find all your angles

OpenStudy (amistre64):

b = 1+c (1+c) +c = sqrt(3) 1 + 2c = sqrt(3) 2c = sqrt(3)-1 c = (sqrt(3)-1)/2; b = (1 + sqrt(3))/2

OpenStudy (anonymous):

I found a=6 then I found , 12/sqrt(3)=(sqrt(3)-1)/sin(B) But couldn't go further

OpenStudy (amistre64):

if your a=6 is good; then: sinA sinB) ---- = ----- a b sinB = b sin(A)/a

OpenStudy (amistre64):

sinB = [ (1 + sqrt(3))/2 ] [sqrt(3)/2] / 6 sinB = (sqrt(3) +3)/24

OpenStudy (amistre64):

sinB = sqrt(3)/24 + 1/8

OpenStudy (amistre64):

the sin inverse funtion will give you one angle for that taio; but you have to be aware that 2 angles are possible with a positive sin.

OpenStudy (amistre64):

taio means ratio..... somehow :)

OpenStudy (anonymous):

Prof ,My problem is how can i turn this to degree. :)

OpenStudy (amistre64):

with the sin inverse function on a calculator.

OpenStudy (anonymous):

I got it , sin inverse function , i will try thanks :)